The weight of students follows a Normal distribution with a mean weight of 60 kg and a standard deviation of 9 kg. 69% of students weigh more than W kg. Find W.
1
Expert's answer
2020-01-06T05:35:16-0500

For a random variable,

"X \sim N(\mu,\sigma^2), \ z=\frac{X-\mu}{\sigma} \sim N(0,1) \newline"

The given weight,


"W \sim N(60,9^2), Z=\frac{W-60}{9} \sim N(0,1) \newline
P(Z>a)=0.69"


From the normal distribution Z table,

"a=-0.50,"


"P(Z>-0.5)=P(\frac{X-\mu}{\sigma}>\frac{W-60}{9}) \newline \because 
W=-0.5 \cdot 9+60=55.5"